# A Novel Quantum Realization of Jet Clustering in High-Energy Physics Experiments

### v2

@@ Acknowledgment
This work is supported by the Fundamental Research Funds for Central Universities, Peking University, the National Key R&D Program of China under Contracts No. 2022YFE0116900, the National Natural Science Foundation of China (NSFC) under Grant No. 12305010, and the Beijing Natural Science Foundation under Grant No. Z220002.
We thank Hongze Xu for his invaluable support in utilizing Quafu Cloud. We extend our appreciation to the Beijing Academy of Quantum Information Sciences for access to quantum hardware and to the Computing Center at the Institute of High-Energy Physics for providing computing resources.
@@
Yongfeng Zhu analyzed the data and wrote the initial manuscript. Weifeng Zhuang, Chen Qian and Yunheng Ma researched QAOA and discussed it with Yongfeng Zhu. Dong E. Liu, Manqi Ruan, and Chen Zhou conceived and oversaw the project. All authors contributed to analyzing the results and to revising and reviewing the manuscript.

### v3

@@ Acknowledgment
This work is supported by the Fundamental Research Funds for Central Universities, Peking University, the state Key Laboratory of Nuclear Physics and Technology, Peking University (Grant No. NPT2022ZZ05), the National Key R&D Program of China under Contracts No. 2022YFE0116900, the National Natural Science Foundation of China (NSFC) under Grant No. 12305010, the Key Laboratory for Particle Astrophysics and Cosmology (Ministry of Education), Shanghai Key Laboratory for Particle Physics and Cosmology, and the Beijing Natural Science Foundation under Grant No. Z220002.
We thank Hongze Xu for his invaluable support in utilizing Quafu Cloud and Andrew Levin for his meticulous efforts in refining the grammar and expressions in our manuscript. We extend our appreciation to the Beijing Academy of Quantum Information Sciences for access to quantum hardware and to the Computing Center at the Institute of High-Energy Physics for providing computing resources.
@@
Yongfeng Zhu analyzed the data and wrote the initial manuscript. Weifeng Zhuang, Chen Qian and Yunheng Ma researched QAOA and discussed it with Yongfeng Zhu. Dong E. Liu, Manqi Ruan, and Chen Zhou conceived of and oversaw the project. All authors contributed to analyzing the results and to revising and reviewing the manuscript.